Grade 12th passPhysical Chemistry

If the velocity of hydrogen molecule is5*104 cm sec1then it's de broglBr wave length is

Sree
7 Years agoGrade 12th pass

1 Answer

Vikas TU
7 Years ago
Hiii Sree
 V of H molecule = 5 *104
Mass Of one mole of H = 2g
Mass of H molecule = 2g/ Avogadro’s no.
= 3.32*10^27 kg
Using De brolie equation
wavelength = h/mv
Wavlength = 4 A0